# RV-Syn: Rational and Verifiable Mathematical Reasoning Data Synthesis Based on Structured Function Library

### v2

@@ 1 Introduction
The development of advanced reasoning Large Language Models (LLMs) Zhao et al. (2023); OpenAI (2024) has markedly improved their ability to address complex tasks across domains such as mathematics, science, and coding. This highlights the importance of synthesizing complex reasoning data to drive further advancements, given the limited availability of high-quality annotated instructions Shah et al. (2024); Yuan et al. (2023).
To address this scarcity, researchers have explored various synthesizing methods, particularly in the mathematics domain. The mainstream methods involve data augmentation based on existing annotated training sets, such as GSM8K Cobbe et al. (2021) and MATH Hendrycks et al. (2021), ranging from self-evolving instructions Xu et al. (2024a); Zeng et al. (2024) and question paraphrasing Yu et al. (2024), to solution augmentation Lu et al. (2024). However, these methods are limited by the available training data, constraining the synthesis diversity Li et al. (2024c, 2023). To enhance diversity, recent approaches enable LLMs to generate a large scale of questions from various mathematics-related sources, including web pages Yue et al. (2024) and knowledge points Tang et al. (2024) from web corpora or textbooks.
@@ 2 Related Work
Researchers have proposed various approaches to enhance the mathematical reasoning capabilities, including methods applied in training or inference stages. During the training stage, existing methods aim to enhance the LLMs from the aspects of pre-training and post-training. Specifically, some studies Azerbayev et al. (2024); Shao et al. (2024); Chen et al. (2024a) involve collecting extensive math-related corpora and enhancing the foundational mathematical capabilities of LLMs through continual pre-training. In contrast, other studies Chen et al. (2024c); Tong et al. (2024); Huang et al. (2024b); Toshniwal et al. (2024) focus on synthesizing a substantial amount of high-quality math-related instructional data, further refining the problem-solving abilities of LLMs through post-training. During the inference stage, there are two primary prompting approaches, including Chain-of-Thought (CoT) Wei et al. (2022) and Program-of-Thought (PoT) Chen et al. (2023). Specifically, some studies Yao et al. (2023); Shinn et al. (2023); Zhou et al. (2023) utilize CoT to elicit LLMs’ inherent reasoning ability with a step-by-step reasoning process. Furthermore, other studies Gou et al. (2024); Wang et al. (2024); Chen et al. (2024b); Wen et al. (2024) incorporate PoT to allow LLMs to utilize external computers during reasoning. In this work, we focus on the training stage and try to synthesize instruction data by leveraging the strengths of both the CoT and PoT methods.
@@
Existing research has proposed various synthesizing approaches: Firstly, existing studies focus on synthesizing new mathematical problems based on few-shot problems Li et al. (2024a), math-related documents Yue et al. (2024), key knowledge points Huang et al. (2024a), or self-evolution Luo et al. (2023). Furthermore, with the improvement of LLMs’ capabilities, subsequent studies Zhou et al. (2024); Ding et al. (2024) utilizingexplore the utilization of open-source LLMs (i.e., DeepSeek-Math and Qwen-Math) to scale the number of synthesized problems. Despite their effectiveness, direct problem generation suffers from a lack of fine-grained control over the problem’s internal logic, leading to lower-quality problems. Besides, these methods frequently lack mechanisms for validating the correctness of generated answers, which may introduce errors and affect the final performance. In contrast, our method starts from the internal logic of problem-solving, rather than directly synthesizing problems. This provides more refined logical consistency and improves problem quality, while also enabling effective correctness control.
@@ 3.4 Comparison to Previous Work
We give a comparison in Table 1. The first line of research, including MetaMath Yu et al. (2024), Orca-Math Mitra et al. (2024), WizardMath Luo et al. (2023), MuggleMath Li et al. (2024b), MathGenie Lu et al. (2024), primarily focuses on augmenting existing problems or solutions, which leads to new instructions that are too similar to the original ones, thus limiting diversity. Another line of research, represented by Mammoth2 Yue et al. (2024), Jiuzhang3.0 Zhou et al. (2024), KPMath Huang et al. (2024a), and ScaleQuest Ding et al. (2024), synthesizes brand new mathematical problems based on math-related documents, knowledge points, or from scratch, offering greater diversity and scalability. However, many of these methods overlook the internal logical processes when directly synthesizing problems. Methods like ControlMath Chen et al. (2024c) and PromptCoT Zhao et al. (2025) attempt to enhance the rationality of synthesis by leveraging equation generators or injecting intermediate thought processes. Nevertheless, they still lack mechanisms to verify the correctness of the synthesized data. Our RV-Syn method approaches problem generation from the perspective of internal problem-solving logic rather than directly producing problem statements without methodical consideration, thereby enhancing problem quality. Additionally, through executable computation graphs, our method enables automatic verification of solution correctness.
@@ 4.1 Experimental Setup
We compare our method with previous problem synthesis methods with publicly available datasets, including: (1) MetaMath Yu et al. (2024) introduces several question bootstrapping techniques; (2) Orca-Math Mitra et al. (2024) augments existing datasets using an Agent-Instruct method; (3) MathScale Tang et al. (2024) uses topic and knowledge-point graphs to prompt new problem synthesis; (4) Mammoth2 Yue et al. (2024) extracts QA pairs from webpages; (5) Jiuzhang3.0 Zhou et al. (2024) uses math-related seed data to synthesize new QA pairs; (6) PromptCoT Zhao et al. (2025) generate complex problems with intermediate thought; (7) ScaleQuest Ding et al. (2024) trains a problem generation model similar to Magpie Xu et al. (2024b).
@@
Directly comparing the performance of models from previous works introduces potential unfair factors, as newer models may leverage more advanced models for backbone and answer generation (see Appendix  F for details). To ensure a fair comparison, we use the same model, Qwen-2.5-Math-7B-Instruct Yang et al. (2024) to generate answers for problems synthesized by each method and train the same backbone model. By controlling for other variables, we can focus on comparing the quality of the synthesized problems.
@@
We assess the models’ performance on MATH-500 Hendrycks et al. (2021); Lightman et al. (2024), GSM8K Cobbe et al. (2021), GSM-Hard Gao et al. (2023), College Math Tang et al. (2024) and OlympiadBench He et al. (2024). The generated outputs are all in the form of natural language Chain-of-Thought (CoT) Wei et al. (2022) through greedy decoding, and we report zero-shot pass@1 accuracy.
@@
We extract seed data from the high-quality Numina-Math dataset LI et al. (2024) to collect function nodes and fine-tuning data. The problem decomposition is performed using Qwen2.5-72B-Instruct. We then use the fine-tuned Qwen2.5-7B-Instruct to generate computational graphs with the strategy described in 3.3.1, followed by problem back-translation using Qwen2.5-72B-Instruct.
@@ 4.3 Scalability Study
In this section, we study the scalability of our method. We compare our method with previous approaches by training on scaling datasets up to 300k using LLaMA-3-8B-Instruct AI@Meta (2024). The results in Figure 3 demonstrate a strong scaling curve, with performance consistently improving as the dataset size increases. Our method consistently outperforms other methods, and notably, achieves better performance with only half the dataset size (50k vs. 100k). This advantage is particularly pronounced on challenging or computation-intensive datasets, demonstrating the superiority of our method for enhancing advanced mathematical reasoning capabilities.
@@ 4.4 Comparison with Human-Crafted Data
To rigorously benchmark the quality of our synthesized data, we compare RV-Syn against datasets that contain a significant portion of high-quality, human-crafted problems. We include two such strong baselines in our analysis: MMIQC Liu et al. (2024), which contains QA pairs from the Mathematics Stack Exchange, and Numina-Math LI et al. (2024), a large collection of both real-world human-crafted and synthesized math problems, which also serves as the source of our initial seed data. As illustrated in Figure 4, human-crafted datasets such as MMIQC and Numina-Math consistently outperform most synthetic methods, reflecting the superior quality of human-authored data over synthetic data, though they face significant scalability challenges. In contrast, RV-Syn demonstrates a clear competitive advantage: at nearly every data scale, it achieves the highest average performance, even surpassing Numina-Math, the very dataset used as its seed. This result validates that our synthesis process does not merely replicate existing examples, but instead generates novel, high-quality instructions that are more effective for model training than the original seed data itself. Consequently, RV-Syn successfully bridges the gap between scalability and quality, producing synthetic data that rivals human expertise without sacrificing performance. We provide further discussion on data diversity in Appendix D.
@@ 4.5 Analysis of Data Efficiency
First, we posit that the solution-first, rational synthesis approach leads to higher-quality and more complex problems. A higher-quality dataset can provide richer supervision signals per example, thus improving training efficiency. To quantify this, we analyze the difficulty and efficacy of different datasets using three key metrics, following PromptCoT Zhao et al. (2025). First, accuracy serves as an indicator of problem difficulty. Lower performance signifies a higher level of difficulty. For datasets without ground-truth labels, we adopt the approach detailed in PromptCoT Zhao et al. (2025) by leveraging a more powerful model, Qwen2.5-Math-72B-Instruct, to generate reference answers, applying self-consistency with 8 rollouts to enhance reliability. Second, the average number of reasoning tokens provides insights into the complexity of the reasoning process. A larger average number of tokens suggests that the model necessitates more extensive reasoning steps. Finally, the average accuracy on the benchmark after tuning reflects the dataset’s contribution to improving model performance. Larger improvements in benchmark accuracy observed after fine-tuning suggest that the respective dataset is more effective in enhancing the model’s capabilities. Experimental results in Table 3 show a strong consistency among these three metrics, and our method achieves the best results across all three difficulty evaluation metrics. Specifically, problems synthesized by RV-Syn are the most difficult (lowest accuracy at 55.6%), require the longest reasoning chains (highest token count), and yield the largest performance improvement after fine-tuning (+34.1%). This demonstrates that our solution-first, rational synthesis approach successfully generates more challenging and effective training instances compared to other methods.
@@ 4.7 Analysis of Long-CoT Distillation
In addition to short-CoT, we also explore the ability of our data to distill Long-CoT abilities. We replace the answer generation model with DeepSeek-R1-Distill-Qwen-7B Guo et al. (2025) and train Qwen2.5-7B-Instruct on a 20k dataset, comparing our method with ScaleQuest. During the evaluation, we set the maximum sequence length to 16k and exclude samples where both models exceed the maximal length. We show the results in Figure 5. We can see that our synthesis method not only performs well on short-CoT but also excels in distilling Long-CoT compared to the baseline.
@@ Appendix D Analysis of Data Diversity
A key limitation of data augmentation methods like MetaMath Yu et al. (2024) is their tendency to produce problems that are structurally similar to the original seeds. This often results from the inherent paraphrasing nature of in-context learning, leading to limited diversity. For example, a seed problem asking to “choose 3 captains from a team of 11 people” might be augmented to “choose 2 captains from a team of 9 people”, a trivial variation. RV-Syn also utilizes seed data, but in a fundamentally different manner. Instead of rewriting whole problems, RV-Syn decomposes them to extract a library of atomic mathematical operations (functions). Our synthesis process then generates new problems not by modifying an existing problem, but by creating entirely new computational graphs by combining these atomic functions. This combinatorial approach allows for the creation of novel reasoning paths and complex problem structures that can be fundamentally different from any single seed problem, thereby fostering genuine diversity.

### v3

@@ 1 Introduction
The development of advanced reasoning Large Language Models (LLMs) survey; openai2024reasoning has markedly improved their ability to address complex tasks across domains such as mathematics, science, and coding. This highlights the importance of synthesizing complex reasoning data to drive further advancements, given the limited availability of high-quality annotated instructions AI-Assisted; scaling.
To address this scarcity, researchers have explored various synthesizing methods, particularly in the mathematics domain. The mainstream methods involve data augmentation based on existing annotated training sets, such as GSM8K gsm8k and MATH dan2021math, ranging from self-evolving instructions wizardlm; Automatic_Instruction_Evolving and question paraphrasing metamath, to solution augmentation mathgenie. However, these methods are limited by the available training data, constraining the synthesis diversity scratch; DBLP:journals/corr/abs-2310-05506. To enhance diversity, recent approaches enable LLMs to generate a large scale of questions from various mathematics-related sources, including web pages mammoth2 and knowledge points mathscale from web corpora or textbooks.
@@ 2 Related Work
Researchers have proposed various approaches to enhance the mathematical reasoning capabilities, including methods applied in training or inference stages. During the training stage, existing methods aim to enhance the LLMs from the aspects of pre-training and post-training. Specifically, some studies llemma; Shao2024DeepSeekMath; syne involve collecting extensive math-related corpora and enhancing the foundational mathematical capabilities of LLMs through continual pre-training. In contrast, other studies DBLP:conf/emnlp/ChenWCS024; dart; MUSTARD; openmathinstruct focus on synthesizing a substantial amount of high-quality math-related instructional data, further refining the problem-solving abilities of LLMs through post-training. During the inference stage, there are two primary prompting approaches, including Chain-of-Thought (CoT) wei2023chain and Program-of-Thought (PoT) pot. Specifically, some studies react; reflexion; least_to_most utilize CoT to elicit LLMs’ inherent reasoning ability with a step-by-step reasoning process. Furthermore, other studies tora; mathcoder; DBLP:journals/corr/abs-2401-05384; codeplan incorporate PoT to allow LLMs to utilize external computers during reasoning. In this work, we focus on the training stage and try to synthesize instruction data by leveraging the strengths of both the CoT and PoT methods.
@@
Existing research has proposed various synthesizing approaches: Firstly, existing studies focus on synthesizing new mathematical problems based on few-shot problems xwin, math-related documents mammoth2, key knowledge points kpmath, or self-evolution wizardmath. Furthermore, with the improvement of LLMs’ capabilities, subsequent studies jiuzhang; scalequest utilizingexplore the utilization of open-source LLMs (i.e., DeepSeek-Math and Qwen-Math) to scale the number of synthesized problems. Despite their effectiveness, direct problem generation suffers from a lack of fine-grained control over the problem’s internal logic, leading to lower-quality problems. Besides, these methods frequently lack mechanisms for validating the correctness of generated answers, which may introduce errors and affect the final performance. In contrast, our method starts from the internal logic of problem-solving, rather than directly synthesizing problems. This provides more refined logical consistency and improves problem quality, while also enabling effective correctness control.
@@ 3.4 Comparison to Previous Work
We give a comparison in Table 1. The first line of research, including MetaMath metamath, Orca-Math orcamath, WizardMath wizardmath, MuggleMath mugglemath, MathGenie mathgenie, primarily focuses on augmenting existing problems or solutions, which leads to new instructions that are too similar to the original ones, thus limiting diversity. Another line of research, represented by Mammoth2 mammoth2, Jiuzhang3.0 jiuzhang, KPMath kpmath, and ScaleQuest scalequest, synthesizes brand new mathematical problems based on math-related documents, knowledge points, or from scratch, offering greater diversity and scalability. However, many of these methods overlook the internal logical processes when directly synthesizing problems. Methods like ControlMath DBLP:conf/emnlp/ChenWCS024 and PromptCoT promptcot attempt to enhance the rationality of synthesis by leveraging equation generators or injecting intermediate thought processes. Nevertheless, they still lack mechanisms to verify the correctness of the synthesized data. Our RV-Syn method approaches problem generation from the perspective of internal problem-solving logic rather than directly producing problem statements without methodical consideration, thereby enhancing problem quality. Additionally, through executable computation graphs, our method enables automatic verification of solution correctness.
@@ 4.1 Experimental Setup
We compare our method with previous problem synthesis methods with publicly available datasets, including: (1) MetaMath metamath introduces several question bootstrapping techniques; (2) Orca-Math orcamath augments existing datasets using an Agent-Instruct method; (3) MathScale mathscale uses topic and knowledge-point graphs to prompt new problem synthesis; (4) Mammoth2 mammoth2 extracts QA pairs from webpages; (5) Jiuzhang3.0 jiuzhang uses math-related seed data to synthesize new QA pairs; (6) PromptCoT promptcot generate complex problems with intermediate thought; (7) ScaleQuest scalequest trains a problem generation model similar to Magpie magpie.
@@
Directly comparing the performance of models from previous works introduces potential unfair factors, as newer models may leverage more advanced models for backbone and answer generation (see Appendix  F for details). To ensure a fair comparison, we use the same model, Qwen-2.5-Math-7B-Instruct Yang2024qwen2.5math to generate answers for problems synthesized by each method and train the same backbone model. By controlling for other variables, we can focus on comparing the quality of the synthesized problems.
@@
We assess the models’ performance on MATH-500 dan2021math; lightman2024step-verify, GSM8K gsm8k, GSM-Hard GaoMZ00YCN23pal, College Math mathscale and OlympiadBench HeLBHTSHHHZLQL024olympiad. The generated outputs are all in the form of natural language Chain-of-Thought (CoT) wei2023chain through greedy decoding, and we report zero-shot pass@1 accuracy.
@@
We extract seed data from the high-quality Numina-Math dataset numina_math_datasets to collect function nodes and fine-tuning data. The problem decomposition is performed using Qwen2.5-72B-Instruct. We then use the fine-tuned Qwen2.5-7B-Instruct to generate computational graphs with the strategy described in 3.3.1, followed by problem back-translation using Qwen2.5-72B-Instruct.
@@ 4.3 Scalability Study
In this section, we study the scalability of our method. We compare our method with previous approaches by training on scaling datasets up to 300k using LLaMA-3-8B-Instruct llama3. The results in Figure 3 demonstrate a strong scaling curve, with performance consistently improving as the dataset size increases. Our method consistently outperforms other methods, and notably, achieves better performance with only half the dataset size (50k vs. 100k). This advantage is particularly pronounced on challenging or computation-intensive datasets, demonstrating the superiority of our method for enhancing advanced mathematical reasoning capabilities.
@@ 4.4 Comparison with Human-Crafted Data
To rigorously benchmark the quality of our synthesized data, we compare RV-Syn against datasets that contain a significant portion of high-quality, human-crafted problems. We include two such strong baselines in our analysis: MMIQC mmiqc, which contains QA pairs from the Mathematics Stack Exchange, and Numina-Math numina_math_datasets, a large collection of both real-world human-crafted and synthesized math problems, which also serves as the source of our initial seed data. As illustrated in Figure 4, human-crafted datasets such as MMIQC and Numina-Math consistently outperform most synthetic methods, reflecting the superior quality of human-authored data over synthetic data, though they face significant scalability challenges. In contrast, RV-Syn demonstrates a clear competitive advantage: at nearly every data scale, it achieves the highest average performance, even surpassing Numina-Math, the very dataset used as its seed. This result validates that our synthesis process does not merely replicate existing examples, but instead generates novel, high-quality instructions that are more effective for model training than the original seed data itself. Consequently, RV-Syn successfully bridges the gap between scalability and quality, producing synthetic data that rivals human expertise without sacrificing performance. We provide further discussion on data diversity in Appendix D.
@@ 4.5 Analysis of Data Efficiency
First, we posit that the solution-first, rational synthesis approach leads to higher-quality and more complex problems. A higher-quality dataset can provide richer supervision signals per example, thus improving training efficiency. To quantify this, we analyze the difficulty and efficacy of different datasets using three key metrics, following PromptCoT promptcot. First, accuracy serves as an indicator of problem difficulty. Lower performance signifies a higher level of difficulty. For datasets without ground-truth labels, we adopt the approach detailed in PromptCoT promptcot by leveraging a more powerful model, Qwen2.5-Math-72B-Instruct, to generate reference answers, applying self-consistency with 8 rollouts to enhance reliability. Second, the average number of reasoning tokens provides insights into the complexity of the reasoning process. A larger average number of tokens suggests that the model necessitates more extensive reasoning steps. Finally, the average accuracy on the benchmark after tuning reflects the dataset’s contribution to improving model performance. Larger improvements in benchmark accuracy observed after fine-tuning suggest that the respective dataset is more effective in enhancing the model’s capabilities. Experimental results in Table 3 show a strong consistency among these three metrics, and our method achieves the best results across all three difficulty evaluation metrics. Specifically, problems synthesized by RV-Syn are the most difficult (lowest accuracy at 55.6%), require the longest reasoning chains (highest token count), and yield the largest performance improvement after fine-tuning (+34.1%). This demonstrates that our solution-first, rational synthesis approach successfully generates more challenging and effective training instances compared to other methods.
@@ 4.7 Analysis of Long-CoT Distillation
In addition to short-CoT, we also explore the ability of our data to distill Long-CoT abilities. We replace the answer generation model with DeepSeek-R1-Distill-Qwen-7B r1 and train Qwen2.5-7B-Instruct on a 20k dataset, comparing our method with ScaleQuest. During the evaluation, we set the maximum sequence length to 16k and exclude samples where both models exceed the maximal length. We show the results in Figure 5. We can see that our synthesis method not only performs well on short-CoT but also excels in distilling Long-CoT compared to the baseline.
@@ Appendix D Analysis of Data Diversity
A key limitation of data augmentation methods like MetaMath metamath is their tendency to produce problems that are structurally similar to the original seeds. This often results from the inherent paraphrasing nature of in-context learning, leading to limited diversity. For example, a seed problem asking to “choose 3 captains from a team of 11 people” might be augmented to “choose 2 captains from a team of 9 people”, a trivial variation. RV-Syn also utilizes seed data, but in a fundamentally different manner. Instead of rewriting whole problems, RV-Syn decomposes them to extract a library of atomic mathematical operations (functions). Our synthesis process then generates new problems not by modifying an existing problem, but by creating entirely new computational graphs by combining these atomic functions. This combinatorial approach allows for the creation of novel reasoning paths and complex problem structures that can be fundamentally different from any single seed problem, thereby fostering genuine diversity.